OpenStudy (akonkel):

A store allows customers to fill their own bags of candy. Glen decides he only wants caramels and candy corn. Caramels sell for $0.95 per pound, and candy corn sells for $0.79 per pound. Glen’s bag weighs 1.7 pounds and it costs $1.48. How many pounds of caramels and candy corn did Glen buy? 0.8 pounds of caramels; 0.9 pounds of candy corn 0.9 pounds of caramels; 0.8 pounds of candy corn 1 pound of caramels; 0.7 pounds of candy corn 0.7 pounds of caramels; 1 pound of candy corn

OpenStudy (studygurl14):

x = caramels y = candy corn 0.95x + 0.79y = 1.48 x + y = 1.7

OpenStudy (akonkel):

HIIII study girl114 thanx i have a few more q's

OpenStudy (anonymous):

can you help me after you are done? i need help on 4 questions and they are worth 25 points

OpenStudy (mathstudent55):

Since you are given choices, you can take each choice, multiply the weight of each candy by its unit price. Add the two prices together, and find which will give you the total the problem tells you.

OpenStudy (studygurl14):

x = 1.7 - y 0.95(1.7 - y) + 0.79y = 1.48 1.615 - 0.95y + 0.79y = 1.48 -0.16y = -0.135 y = 0.84375 pountds, rounded to 0.8

OpenStudy (studygurl14):

x = 1.7 - y x = 1.7 - (0.8) x = 0.9

OpenStudy (akonkel):

alright i got it now thanx guys (all) i just have to more q's

OpenStudy (akonkel):

ready?